# On Polar Coding with Feedback

### v1

@@ IV Simulation Results
The block error rate (BLER) performance of our proposed feedback coding scheme for BSC(0.11) is shown in Fig. 6. For comparison, the performance curves without feedback are also provided to demonstrate the improvement. When $D_{\max}=\infty$, the system attains theoretical error-free performance. This point is marked at a BLER of $10^{-6}$ in the figure. For other points, we adjust the threshold $\epsilon_{\operatorname{\mathrm{th}}}$ as in Tabel I to control the distribution of $|\mathcal{T}|$ and $D$ for different tolerance $D_{\max}$. These points characterize the trade-off between rate and feedback delay. Thanks to Theorem 2, $P(|\mathcal{T}|)$ can be accurately estimated. which, in turn, enables precise prediction of the BLER.
@@
As another by-product, we show that Theorem 2 can be employed to compress the number $|\mathcal{T}|$, when the polar lossless coding scheme [19] is operating in a serial data-stream mode, where $|\mathcal{T}|$ is required for truncation. For different thresholds, we derive an empirical estimate of $P(|\mathcal{T}|)$ via Monte Carlo simulation. The entropy of $P(|\mathcal{T}|)$ is denoted by $H(|\mathcal{T}|)$. We then use $P_{\mathrm{NB}}(|\mathcal{T}|)$ from Theorem 2, with entropy $H_{\mathrm{NB}}(|\mathcal{T}|)$, to generate a Huffman dictionary for $|\mathcal{T}|$ and compute its average length $\bar{L}$. $H(|\mathcal{T}|)$, $H_{\mathrm{NB}}(|\mathcal{T}|)$ and $\bar{L}$ (all in bits) are listed in Tabel II, which shows the validity of our statistical model.
TABLE II: The compression of $|\mathcal{T}|$ using the distribution in Theorem 2.

### v2

@@ IV Simulation Results
The block error rate (BLER) performance of our proposed feedback coding scheme for BSC(0.11) is shown in Fig. 6. For comparison, the performance curves without feedback are also provided to demonstrate the improvement. When $D_{\max}=\infty$, the system attains theoretical error-free performance. This point is marked at a BLER of $10^{-6}$ in the figure. For other points, we adjust the threshold $\epsilon_{\operatorname{\mathrm{th}}}$ as in Table I to control the distribution of $|\mathcal{T}|$ and $D$ for different tolerance $D_{\max}$. These points characterize the trade-off between rate and feedback delay. Thanks to Theorem 2, $P(|\mathcal{T}|)$ can be accurately estimated, which, in turn, enables precise prediction of the BLER.
@@
As another by-product, we show that Theorem 2 can be employed to compress the number $|\mathcal{T}|$, when the polar lossless coding scheme [19] is operating in a serial data-stream mode, where $|\mathcal{T}|$ is required for truncation. For the i.i.d. Bernoulli (BER) source with probability 0.11 and length $N=2^{10}$, we derive an empirical estimate of $P(|\mathcal{T}|)$ via Monte Carlo simulation for different thresholds. The entropy of $P(|\mathcal{T}|)$ is denoted by $H(|\mathcal{T}|)$. We then use $P_{\mathrm{NB}}(|\mathcal{T}|)$ from Theorem 2, with entropy $H_{\mathrm{NB}}(|\mathcal{T}|)$, to generate a Huffman dictionary for $|\mathcal{T}|$ and compute its average length $\bar{L}$. $H(|\mathcal{T}|)$, $H_{\mathrm{NB}}(|\mathcal{T}|)$ and $\bar{L}$ (all in bits) are listed in Table II, which shows the validity of our statistical model.
TABLE II: The compression of $|\mathcal{T}|$ for Ber(0.11) and $N=2^{10}$, using the distribution in Theorem 2.